# DEL-Ranking: Ranking-Correction Denoising Framework for Elucidating Molecular Affinities in DNA-Encoded Libraries

### v1

@@ 3.3 Activity-Referenced Distribution Correction Framework
To effectively leverage activity information and address distribution shifts in DEL experiments, we propose the Activity-Referenced Distribution Correction (ARDC) framework. This approach reinforces the impact of activity labels on read count distributions through two key components: Iterative Activity-Guided Refinement (AGR) and Consistency-Driven Error Correction (CDEC).
The Iterative Activity-Guided Refinement process incorporates activity information into read count predictions. Detailed in Algorithm 1, this process handles matrix counts and target counts separately. Matrix counts are predicted using 2D SMILES embeddings, while target counts leverage 2D-3D joint embeddings. An adaptive iterative mechanism, inspired by self-training techniques in semi-supervised learning (Zoph et al., 2020), progressively improves the consistency between read count predictions and activity labels. This recursive updating approach reinforces the influence of activity labels on read count values, ultimately converging towards more biologically consistent results.
To address potential error accumulation and further align predictions with biological reality, we introduce the Consistency-Driven Error Correction component. This feature utilizes a consistency loss function that not only regresses the predicted values but also aligns their trends with activity labels. This approach helps resolve discrepancies between compounds with low read counts but high activity. The consistency loss is defined as:
$\mathcal{L}_{\text{consist}}=\|\mathbf{A}_{\text{pred}}-\mathbf{A}_{\text{true}}\|+\max\Big{(}0,\|\mathbf{A}_{\text{pred}}-\hat{\mathbf{R}}_{\text{pred}}\|_{2}^{2}-\|\mathbf{A}_{\text{true}}-\hat{\mathbf{R}}_{\text{true}}\|_{2}^{2}\Big{)}$ (8)
where $\mathbf{A}_{\text{pred}}$ and $\mathbf{A}_{\text{true}}$ denote the predicted and ground-truth activity, respectively, while $\hat{\mathbf{R}}_{\text{pred}}$ and $\hat{\mathbf{R}}_{\text{true}}$ represent the normalized predicted and ground-truth read count values between 0 and 1.
@@ 3.4 Total Training Objective
The total training objective integrates three distinct components. The zero-inflated distribution loss $\mathcal{L}_{\text{ZIP}}$ models the overall read count distribution, while the combined ranking loss $\mathcal{L}_{\text{rank}}$ refines the predicted ZIP distribution based on ordinal relationships. Additionally, the consistency loss $\mathcal{L}_{\text{consist}}$ further adjusts the distribution using activity labels. These components are combined into the total loss function as follows:
@@ 4 Experiment
Datasets. CA9 Dataset: From the original data containing 108,529 DNA-barcoded molecules targeting human carbonic anhydrase IX (CA9) (Gerry et al., 2019), we derived two separate datasets. The first, denoted as 5fl4-9p, uses 9 docked poses that we generated ourselves. The second, 5fl4-20p, employs 20 docked poses using the 5fl4 structure. Both datasets lack activity labels. CA2 and CA12 Datasets From the CAS-DEL library (Hou et al., 2023), we generated three datasets comprising 78,390 molecules selected from 7,721,415 3-cycle peptide compounds. We performed docking to create 9 poses per molecule for each dataset. The CA2-derived dataset uses the 3p3h PDB structure (denoted as 3p3h), while two CA12-derived datasets use the 4kp5 PDB structure: 4kp5-A for normal expression and 4kp5-OA for overexpression. Binary affinity labels (0/1) are assigned based on a specific building block (BB3-197). Validation Dataset from ChEMBL (Zdrazil et al., 2024) includes 12,409 small molecules with affinity measurements for CA9, CA2, and CA12. Molecules have compatible atom types, molecular weights from 25 to 1000 amu, and inhibitory constants (Ki) from 90 pM to 0.15 M. A subset focusing on the 10-90th percentile range of the training data’s molecular weights provides a more challenging test scenario.
Evaluation Metrics and Hyperparameters. To evaluate our framework’s effectiveness, we employ two Spearman correlation metrics on the ChEMBL dataset (Zdrazil et al., 2024). The first metric, overall Spearman correlation ($\rho_{\text{overall}}$), measures the correlation between predicted read counts and experimentally determined $K_{i}$ values across the entire validation dataset. Also, we utilize the subset Spearman correlation ($\rho_{\text{subset}}$), which focuses on compounds with molecular weights within the 10th to 90th percentile range of the training dataset. Hyperparameters setting is shown in Appendix A.3.
Baselines. We examine the performance of existing binding affinity predictors. Traditional methods based on binding poses and fingerprints inculde Molecule Weight, Benzene Sulfonamide, Vina Docking (Koes et al., 2013), and Dos-DEL(Gerry et al., 2019). AI-aided methods dependent of read count values and molecule information include DEL-QSVR, and DEL-Dock (Lim et al., 2022; Shmilovich et al., 2023). All methods are tested on 5 different target datasets.
@@ 4.1 Performance Comparison
Zero-shot Generalization. We tested zero-shot generalization ability of DEL-Dock (Shmilovich et al., 2023) and DEL-Ranking on CA9 target, trained on 3p3h, 4kp5-A, and 4kp5-OA datasets. As shown in Table 3, DEL-Ranking consistently outperforms DEL-Dock across all datasets, demonstrating the robustness of our approach in diverse conditions. Notably, for the 4kp5-OA dataset, which likely contains higher experimental noise and poses significant challenges, DEL-Ranking shows a substantial improvement, revealing its enhanced robustness in handling noisy DEL data.
@@ 4.2 Discovery of potential high affinity functional group
To assess DEL-Ranking’s ability in identifying high-affinity compounds and exploring potential functional groups, we analyzed Top-50 target count samples predicted by our model on five datasets. As shown in Figure 3, the majority of these samples demonstrated consistently low Ki values, validating the model’s effectiveness in prioritizing active compounds within large DEL libraries.
Known Group Accuracy. DEL-Ranking shows exceptional accuracy in detecting benzene sulfonamide, a key high-affinity group for carbonic anhydrase inhibitors (Hou et al., 2023). From Figure 3, we can observe that the model achieved high detection rates on five datasets, indicating effective incorporation of chemical space information through activity labels.
@@
Further investigation through case-by-case Ki value determination yielded compelling results. Five compounds from 3p3h and five from 5fl4 containing pyrimidine sulfonamide exhibited $K_{i}$ values comparable to or even surpassing those of benzene sulfonamide-containing compounds. This finding profoundly validates DEL-Ranking’s dual capability: successfully incorporating activity label information, while simultaneously leveraging multi-level information along with integrated ranking orders to uncover potential high-activity functional groups. Notably, this discovery reveals DEL-Ranking’s ability to identify unexplored scaffolds, showing potential to improve compound prioritization and accelerate hit-to-lead optimization in early-stage drug discovery. Detailed visualization of Top-50 samples and selected Pyrimidine Sulfonamide cases are shown in  Sections A.4 and A.5.
@@ 4.3 Ablation Study
To further explore the effectiveness of our enhancement, we compare DEL-Ranking with some variants on 3p3h, 4kp5-A, and 4kp5-OA datasets. We can observe from Table 4 that (1) $\mathcal{L}_{\text{PSR}}$ and $\mathcal{L}_{\text{LGR}}$ contribute most significantly to model performance across all datasets. (2) The impact of $\mathcal{L}_{\text{PSR}}$ is more pronounced in datasets with higher noise levels, as evidenced by the larger relative performance drop in the 3p3h dataset. (3) Temperature adjustment and $\mathcal{L}_{c}$ help improve the performance by correcting the predicted distributions, but count less than ranking-based denoising.
@@
w/o PSR -0.273$\pm$ 0.012 -0.155$\pm$ 0.013 -0.251$\pm$ 0.015 -0.271$\pm$ 0.011 0.015$\pm$ 0.028 -0.105$\pm$ 0.033
w/o LGR -0.280$\pm$ 0.011 -0.168$\pm$ 0.015 -0.256$\pm$ 0.023 -0.273$\pm$ 0.016 -0.269$\pm$ 0.024 -0.209$\pm$ 0.034
w/o LGR-con -0.283$\pm$ 0.004 -0.172$\pm$ 0.007 -0.260$\pm$ 0.018 -0.273$\pm$ 0.014 -0.273$\pm$ 0.024 -0.218$\pm$ 0.034
@@ A.1 Proof of Lemma and Theorem
[Proof of Lemma1] Let $(\Omega,\mathcal{F},P)$ be a probability space and $(X,Y):\Omega\to\mathcal{X}\times\mathbb{N}_{0}$ be random variables representing features and read counts respectively. Define $f_{ZIP}(y|x)$ as the probability mass function of a well-fitted Zero-Inflated Poisson model.
@@
$\hat{Y}(x)=E[Y|X=x]$ $=\sum_{y=0}^{\infty}y\cdot f_{ZIP}(y|x)$
$\mathcal{L}_{\text{ZIP}}(f_{ZIP},\mathcal{D})$ $=-\sum_{(x,y)\in\mathcal{D}}\log f_{ZIP}(y|x)$
@@
$\mathcal{L}_{\text{ZIP}}(f_{ZIP},\{(x_{i},y_{i}),(x_{j},y_{j})\})$ $=-\log f_{ZIP}(y_{i}|x_{i})-\log f_{ZIP}(y_{j}|x_{j})$
@@
$E[L_{C}(\hat{y}^{C}(\alpha),Y;\alpha)]\leq E[L_{C}(\hat{y}^{ZI},Y;\alpha)]$
@@
$\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C}(\alpha),Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{C}(\alpha),Y)]\leq\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{ZI},Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{ZI},Y)]$
Let $\Delta\mathcal{L}_{\text{ZIP}}(\alpha)=E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C}(\alpha),Y)]-E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{ZI},Y)]$ and $\Delta\mathcal{L}_{\text{rank}}(\alpha)=E[\mathcal{L}_{\text{rank}}(\hat{y}^{ZI},Y)]-E[\mathcal{L}_{\text{rank}}(\hat{y}^{C}(\alpha),Y)]$. Rearranging the inequality:
@@
$E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C}(\alpha),Y)]-E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{ZI},Y)]<\frac{1-\alpha}{\alpha}(E[\mathcal{L}_{\text{rank}}(\hat{y}^{ZI},Y)]-E[\mathcal{L}_{\text{rank}}(\hat{y}^{C}(\alpha),Y)])$
@@
$E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C},Y)]-E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{ZI},Y)]<\frac{1-\alpha}{\alpha}(E[\mathcal{L}_{\text{rank}}(\hat{y}^{ZI},Y)]-E[\mathcal{L}_{\text{rank}}(\hat{y}^{C},Y)])$
@@
$\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C},Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{C},Y)]<\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{ZI},Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{ZI},Y)]$
The left-hand side of this inequality is $E[L_{C}(\hat{y}^{C})]$ by definition. The right-hand side is strictly greater than $E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{ZI})]$ since $E[\mathcal{L}_{\text{rank}}(\hat{y}^{ZI},Y)]>0$ for any non-trivial ranking loss and $\alpha<1$.
@@
$E[L_{C}(\hat{y}^{C})]<\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{ZI},Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{ZI},Y)]<E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{ZI})]$

### v2

@@ 3.3 Activity-Referenced Distribution Correction Framework
To effectively leverage activity information and address Distribution Shifts in DEL experiments, we propose the Activity-Referenced Correction (ARC) framework. This algorithm enhances read count distribution fitting through two complementary stages: the Refinement Stage, which integrates activity information, and the Consistency Stage, which performs distribution adjustment, jointly optimizing the fitting from both dimensions. The Refinement Stage leverages dual information streams - activity labels and read counts - which reflect compound activity from complementary angles: activity labels capture overall binding potential, while read counts quantify binding strength. As detailed in Algorithm 1, our approach generates initial predictions using both 2D SMILES embeddings and 2D-3D joint embeddings for matrix and target counts. To fuse these predictions effectively, we implement an adaptive iterative mechanism inspired by self-training techniques (Zoph et al., 2020). Through multiple rounds of updates, a bidirectional feedback loop is established: activity information calibrates noisy read count predictions, while read count patterns help validate activity predictions. This iterative refinement ensures biological consistency while enhancing prediction accuracy through mutual correction between both information spaces.
To address error accumulation and better align predictions with biological reality, we introduce a consistency loss function in the Correction Stage. This function not only regresses predicted values but also aligns prediction trends with activity labels. Following  (Hou et al., 2023), we define ground-truth labels based on the presence of benzene sulfonamide in molecules. Although there may exist high-affinity molecules without benzene sulfonamide, this labeling scheme provides effective supervision signals for both read count regression and novel high-affinity functional group discovery, as demonstrated in Section 4.2. This approach helps resolve discrepancies in compounds that exhibit low read counts but high activity. The consistency loss is defined as:
$\mathcal{L}_{\text{consist}}(r_{i},\hat{r}_{i},y_{i},\hat{y}_{i})=\|\hat{y}_{i}-y_{i}\|+\max\Big{(}0,\|\hat{y}_{i}-\frac{\hat{r}_{i}}{\max_{i\in\{1,...,N\}}\hat{r}_{i}}\|_{2}^{2}-\|y_{i}-\frac{r_{i}}{\max_{i\in\{1,...,N\}}r_{i}}\|_{2}^{2}\Big{)}$ (10)
where N denotes training batch size. The first term ensures prediction accuracy, while the second term constrains the consistency between read count values and binary activity predictions.
@@ 3.4 Total Training Objective
The total loss function combines three components: Zero-Inflated Poisson loss $\mathcal{L}_{\text{ZIP}}$ for modeling read count distribution, ranking loss $\mathcal{L}_{\text{rank}}$ for optimizing predictions based on ordinal relationships, and consistency loss $\mathcal{L}_{\text{consist}}$ for adjusting distribution using activity labels.
@@ 4 Experiment
Datasets. CA9 Dataset From the original data containing 108,529 DNA-barcoded molecules targeting human carbonic anhydrase IX (CA9) (Gerry et al., 2019), we derived two separate datasets. The first, denoted as 5fl4-9p, uses 9 docked poses that we generated ourselves. The second, 5fl4-20p, employs 20 docked poses using the 5fl4 structure. Both datasets lack activity labels. CA2 and CA12 Datasets From the CAS-DEL library (Hou et al., 2023), we generated three datasets comprising 78,390 molecules selected from 7,721,415 3-cycle peptide compounds. We performed docking to create 9 poses per molecule for each dataset. The CA2-derived dataset uses the 3p3h PDB structure (denoted as 3p3h), while two CA12-derived datasets use the 4kp5 PDB structure: 4kp5-A for normal expression and 4kp5-OA for overexpression. The binary activity label is set to 1 when there is benzene sulfonamide (BB3-197) in the compound (Hou et al., 2023). Validation Dataset from ChEMBL (Zdrazil et al., 2024) includes 12,409 small molecules with affinity measurements for CA9, CA2, and CA12. Molecules have compatible atom types, molecular weights from 25 to 1000 amu, and inhibitory constants (Ki) from 90 pM to 0.15 M. A subset focusing on the 10-90th percentile range of the training data’s molecular weights provides a more challenging test scenario. Virtual Docking details for ligand poses are shown in Appendix B.1.
Evaluation Metrics and hyper-parameters. To evaluate our framework’s effectiveness, we employ two Spearman correlation metrics on the ChEMBL dataset (Zdrazil et al., 2024). The first metric, overall Spearman correlation ($\rho_{\text{overall}}$), measures the correlation between predicted read counts and experimentally determined $K_{i}$ values across the entire validation dataset. Also, we utilize the subset Spearman correlation ($\rho_{\text{subset}}$), which focuses on compounds with molecular weights within the 10th to 90th percentile range of the training dataset. Since multiple hyper-parameters are shown in DEL-Ranking method, we provide a detailed hyper-parameter ssetting in Appendix B.2.
Baselines. We examine the performance of existing binding affinity predictors. Traditional methods based on binding poses and fingerprints include Molecule Weight, Benzene Sulfonamide, Vina Docking (Koes et al., 2013), and Dos-DEL(Gerry et al., 2019). AI-aided methods dependent on read count values and molecule information include RF-Enrichment, RF-ZIP (Random Forest for Log-enrichment, $\mathcal{L}_{\text{ZIP}}$), DEL-QSVR, and DEL-Dock (Lim et al., 2022; Shmilovich et al., 2023).
@@ 4.1 Performance Comparison
Zero-shot Generalization. We evaluated models’ zero-shot generalization on CA9 by training them on CA2 and CA12 targets across three datasets (3p3h, 4kp5-A, and 4kp5-OA). Detailed in Table 3, DEL-Ranking consistently outperformed DEL-Dock. Notably, on the 4kp5-OA dataset with substantially different protein targets, DEL-Ranking maintained strong predictive performance, demonstrating its generalization capability to novel targets. Interestingly, DEL-QSVR exhibited superior zero-shot performance, suggesting that simpler molecular representations and loss functions might be more conducive to target generalization. This superior performance might be attributed to the fact that incorporating pose information could potentially limit zero-shot generalization capability.
@@ 4.2 Discovery of potential high affinity functional group
To evaluate DEL-Ranking’s capability in identifying potent compounds, we analyzed the Top-50 cases from our model across five datasets. Detailed in Figure 3, selected compounds consistently exhibited low Ki values, revealing the model’s ability to prioritize high-affinity compounds from large DEL libraries.
Known Group Accuracy. DEL-Ranking shows expectational accuracy in detecting benzene sulfonamide, a key high-affinity group for carbonic anhydrase inhibitors (Hou et al., 2023). From Figure 3, the model achieved high detection rates on five datasets, demonstrating that our ARC framework effectively incorporates biological prior knowledge into model prediction. To further explore the potential high-affinity compounds, we conducted the same study of DEL-Dock (Shmilovich et al., 2023) in Appendix C.3.
@@
Further investigation through case-by-case Ki value determination yielded compelling results. Five compounds from 3p3h and five from 5fl4 containing pyrimidine sulfonamide exhibited $K_{i}$ values comparable to or even surpassing those of benzene sulfonamide-containing compounds. This finding profoundly validates DEL-Ranking’s dual capability: successfully incorporating activity label information, while simultaneously leveraging multi-level information along with integrated ranking orders to uncover potential high-activity functional groups. Notably, this discovery reveals DEL-Ranking’s ability to identify unexplored scaffolds, showing potential to improve compound prioritization and accelerate hit-to-lead optimization in early-stage drug discovery. Detailed visualization of Top-50 samples and selected Pyrimidine Sulfonamide cases are shown in  Sections C.3 and C.5.
@@
Mol Weight -0.121 -0.028 -0.121 -0.028 -0.121 -0.028
Benzene -0.174 -0.134 -0.174 -0.134 -0.174 -0.134
Vina Docking -0.114±0.009 -0.055±0.007 -0.114±0.009 -0.055±0.007 -0.114±0.009 -0.055±0.007
RF-Enrichment 0.020±0.014 -0.031±0.057 -0.034±0.013 -0.034±0.029 -0.044±0.005 -0.085±0.006
RF-ZIP 0.037±0.059 0.013±0.017 0.036±0.024 -0.002±0.016 0.049±0.012 -0.007±0.013
Dos-DEL -0.115±0.065 -0.036±0.010 -0.115±0.065 -0.036±0.010 -0.115±0.065 -0.036±0.010
DEL-QSVR -0.300±0.020 -0.257±0.022 -0.236±0.038 -0.223±0.030 0.108±0.089 0.130±0.070
@@ 4.3 Ablation Study
To further explore the effectiveness of our enhancement, we compare DEL-Ranking with some variants on 3p3h, 4kp5-A, and 4kp5-OA datasets. We can observe from Table 4 that (1) $\mathcal{L}_{\text{PSR}}$ and $\mathcal{L}_{\text{LGR}}$ contribute most significantly to model performance across all datasets. (2) The impact of $\mathcal{L}_{\text{PSR}}$ is more pronounced in datasets with higher noise levels, as evidenced by the larger relative performance drop in the 3p3h dataset. (3) Temperature adjustment and $\mathcal{L}_{\text{consist}}$ help improve the performance by correcting the predicted distributions, but count less than ranking-based denoising.
@@
w/o $\mathcal{L}_{\text{PSR}}$0.273$\pm$ 0.012 -0.155$\pm$ 0.013 -0.251$\pm$ 0.015 -0.271$\pm$ 0.011 0.015$\pm$ 0.028 -0.105$\pm$ 0.033
w/o $\mathcal{L}_{\text{LGR}}$0.280$\pm$ 0.011 -0.168$\pm$ 0.015 -0.256$\pm$ 0.023 -0.273$\pm$ 0.016 -0.269$\pm$ 0.024 -0.209$\pm$ 0.034
w/o $\mathcal{L}_{\text{con}}$0.283$\pm$ 0.004 -0.172$\pm$ 0.007 -0.260$\pm$ 0.018 -0.273$\pm$ 0.014 -0.273$\pm$ 0.024 -0.218$\pm$ 0.034
@@
To further validate the robustness of DEL-Ranking and the incorporation of additional pose 3D structure information, we conducted ablation studies on both loss weight and structure information weight. The detailed results can be found in  Sections C.2 and C.1. The experimental results corroborate the capability of our approach and the feasibility of our hyperparameter selection criteria.
@@ A.1 Proof of Lemma and Theorem
[Proof of Lemma1] Let $(\Omega,\mathcal{F},P)$ be a probability space and $(X,Y):\Omega\to\mathcal{X}\times\mathbb{N}_{0}$ be random variables representing features and read counts respectively. Define $f_{\text{ZIP}}(y|x)$ as the probability mass function of a well-fitted Zero-Inflated Poisson model.
@@
$\hat{Y}(x)=E[Y|X=x]$ $=\sum_{y=0}^{\infty}y\cdot f_{\text{ZIP}}(y|x)$
$\mathcal{L}_{\text{ZIP}}(f_{\text{ZIP}},\mathcal{D})$ $=-\sum_{(x,y)\in\mathcal{D}}\log f_{\text{ZIP}}(y|x)$
@@
$\mathcal{L}_{\text{ZIP}}(f_{\text{ZIP}},\{(x_{i},y_{i}),(x_{j},y_{j})\})$ $=-\log f_{\text{ZIP}}(y_{i}|x_{i})-\log f_{\text{ZIP}}(y_{j}|x_{j})$
@@
$E[L_{C}(\hat{y}^{C}(\alpha),Y;\alpha)]\leq E[L_{C}(\hat{y}^{\text{ZIP}},Y;\alpha)]$
@@
$\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C}(\alpha),Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{C}(\alpha),Y)]\leq\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{\text{ZIP}},Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{\text{ZIP}},Y)]$
Let $\Delta\mathcal{L}_{\text{ZIP}}(\alpha)=E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C}(\alpha),Y)]-E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{\text{ZIP}},Y)]$ and $\Delta\mathcal{L}_{\text{rank}}(\alpha)=E[\mathcal{L}_{\text{rank}}(\hat{y}^{\text{ZIP}},Y)]-E[\mathcal{L}_{\text{rank}}(\hat{y}^{C}(\alpha),Y)]$. Rearranging the inequality:
@@
$E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C}(\alpha),Y)]-E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{\text{ZIP}},Y)]<\frac{1-\alpha}{\alpha}(E[\mathcal{L}_{\text{rank}}(\hat{y}^{\text{ZIP}},Y)]-E[\mathcal{L}_{\text{rank}}(\hat{y}^{C}(\alpha),Y)])$
@@
$E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C},Y)]-E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{\text{ZIP}},Y)]<\frac{1-\alpha}{\alpha}(E[\mathcal{L}_{\text{rank}}(\hat{y}^{\text{ZIP}},Y)]-E[\mathcal{L}_{\text{rank}}(\hat{y}^{C},Y)])$
@@
$\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{C},Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{C},Y)]<\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{\text{ZIP}},Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{\text{ZIP}},Y)]$
The left-hand side of this inequality is $E[L_{C}(\hat{y}^{C})]$ by definition. The right-hand side is strictly greater than $E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{\text{ZIP}})]$ since $E[\mathcal{L}_{\text{rank}}(\hat{y}^{\text{ZIP}},Y)]>0$ for any non-trivial ranking loss and $\alpha<1$.
@@
$E[L_{C}(\hat{y}^{C})]<\alpha E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{\text{ZIP}},Y)]+(1-\alpha)E[\mathcal{L}_{\text{rank}}(\hat{y}^{\text{ZIP}},Y)]<E[\mathcal{L}_{\text{ZIP}}(\hat{y}^{\text{ZIP}})]$